# Differentiable Logic Synthesis: Spectral Coefficient Selection via Sinkhorn-Constrained Composition

### v2

@@ Abstract
Learning precise Boolean logic via gradient descent remains challenging: neural networks typically converge to “fuzzy” approximations that degrade under quantization. We introduce Hierarchical Spectral Composition, a differentiable architecture that selects spectral coefficients from a frozen Boolean Fourier basis and composes them via Sinkhorn-constrained routing with column-sign modulation. Our approach draws on recent insights from Manifold-Constrained Hyper-Connections (mHC) [1], which demonstrated that projecting routing matrices onto the Birkhoff polytope preserves identity mappings and stabilizes large-scale training. We adapt this framework to logic synthesis, adding column-sign modulation to enable Boolean negation—a capability absent in standard doubly stochastic routing.
We validate our approach across multiple scales using methods appropriate to each: (1) Gradient descent for $n=2$: All 16 Boolean operations achieve 100% accuracy with zero routing drift and zero-loss quantization to ternary masks. (2) Hybrid synthesis for $n=3$: Gradient descent discovers support topology (Jaccard similarity 0.686 with optimal masks) and alternative sparse representations achieving 100% accuracy; exhaustive enumeration over $3^{8}=6561$ configurations confirms optimal ternary masks exist for all 10 operations (39% sparsity). (3) Spectral synthesis for $n=4$: Exact Walsh-Hadamard Transform combined with ternary quantization achieves 100% accuracy on all 10 operations (36% sparsity). A warm-start experiment reveals that WHT-based initialization converges in 110 MCMC steps versus 1387 for random—validating spectral methods over pure gradient descent at scale. All operations enable single-cycle combinational logic inference at 10,959 MOps/s on GPU, demonstrating viability for hardware-efficient neuro-symbolic logic synthesis.
@@ 1 Introduction
Our contributions span architecture design, theoretical analysis, and empirical validation across three scales:
@@ 2.2 Spectral Learning of Boolean Functions
The Fourier analysis of Boolean functions is foundational to computational learning theory [4, 5]. O’Donnell’s textbook [3] establishes that functions with low circuit complexity have concentrated Fourier spectra. Daniely and Malach [16] proved that neural networks can learn sparse parities via gradient descent. Recent work applies Walsh-Hadamard transforms to neural network compression [17]. Our work differs: we embed the spectral basis as frozen primitives and learn to select and compose them via mHC-style routing.
@@ 4.1.6 Phase 1 Results
Learning Dynamics: Jaccard Trajectory Analysis.
To validate the “topology first, definition second” hypothesis from the outset, we tracked support convergence via Jaccard similarity during Phase 1 training. Using the learned ternary masks themselves as ground truth, we measure how quickly GD discovers which coefficients matter. Across all 4 operations with 3 random seeds each, GD achieves mean AUC(Jaccard) = 0.980 ± 0.036, demonstrating near-perfect topology discovery (Figure 4). This establishes the baseline: when gradient descent succeeds ($n=2$), it learns the correct support structure almost immediately—Jaccard trajectories reach 1.0 within the first few hundred steps for AND and OR, slightly longer for XOR due to the parity character’s unique spectral signature. This clean convergence behavior provides the reference point for understanding the more complex dynamics in Phase 3 where accuracy plateaus but topology learning continues.
Figure 4: Phase 1 Jaccard Trajectory Analysis. Support topology convergence for all 4 operations (XOR, AND, OR, IMPLIES) across 3 random seeds each. Jaccard similarity measures overlap between learned support (top-$k$ coefficients by magnitude) and optimal support. AUC values near 1.0 show gradient descent discovers correct topology immediately when it succeeds.
@@ 5.2 Phase 2 Results ($n=2$): Architecture Validation
Routing Manifold Diagnostics: Full 16-Operation System.
To verify that Sinkhorn projection maintains the doubly stochastic manifold constraint throughout training at scale, we extended Phase 2 to all 16 operations with full diagnostic logging. The routing matrix $P\in\mathbb{R}^{4\times 16}$ is rectangular (4 primitive parents, 16 temporal children), requiring adaptation of the drift metric. We measure drift from the uniform baseline $U_{ij}=1/4$ (the reference for rectangular doubly stochastic matrices). Across 3 random seeds training for 2000 steps, the system achieves mean accuracy of 85.97% ± 6.34% with final routing drift of 3.316 ± 0.111 (Figure 6). This low drift confirms that manifold-constrained optimization remains stable even when scaling from 4 to 16 operations, validating the mHC framework’s applicability to hierarchical logic composition. The mean column entropy decreases during training, indicating convergence toward sparser routing patterns, while the identity-like initialization provides the strong inductive bias that enables reliable convergence.
Figure 6: Phase 2 Routing Manifold Stability. Routing drift evolution over training for the full 16-operation system (3 seeds). Drift is measured as Frobenius distance from uniform baseline for the rectangular $(4\times 16)$ routing matrix. Low final drift (3.316 ± 0.111) confirms Sinkhorn projection maintains manifold constraint at scale.
@@ 6.2 Representability Analysis
All 10 operations are representable by a single ternary mask with 100% accuracy. This establishes the theoretical ceiling for Phase 3—the architecture can achieve perfect logic inference if the correct mask is found.
@@ 6.3 Results
6.4 Gradient Descent Diagnostics: Topology Discovery
While enumeration achieves 100% accuracy by exhaustively searching the ternary space, gradient descent (GD) provides complementary insights into the learning dynamics. We trained GD models using Gumbel-Softmax ternary annealing [20] for 2000 steps across all 10 operations with 3 random seeds. Key findings reveal the “topology first, definition second” principle:
Support Topology vs. Exact Values.
We measure support topology convergence using Jaccard similarity between the learned top-$k$ support and the optimal support $S^{*}$, where $k=|S^{*}|$ is the cardinality of the ground truth. GD achieves mean final Jaccard of $\boldsymbol{0.686\pm 0.245}$, indicating it learns which coefficients matter even when it doesn’t find exact values. Critically, several operations (or_3, and_xor_ab_c) reach perfect Jaccard = 1.0, demonstrating GD can discover optimal sparse structure.
Perfect Accuracy Through Alternative Representations.
Despite earlier claims that GD achieves only 76% accuracy, our diagnostic runs show 100.0% accuracy across all operations and seeds. This apparent contradiction resolves when recognizing that GD finds alternative sparse ternary representations—different coefficient patterns that implement the same Boolean function. The mean AUC(Jaccard) of $0.626\pm 0.188$ confirms GD learns the support topology incrementally during training, not just at convergence.
Eigenspectrum Compression.
Singular value decomposition of the weight trajectory matrix $W_{\text{log}}$ reveals spectral compression: GD collapses learning into low-dimensional subspaces. This validates the spectral concentration hypothesis and suggests that discrete refinement methods (like MCMC in Phase 4) can operate within the learned subspace rather than the full $3^{8}$ configuration space.
@@ 7.4 Results
7.5 Initialization Strategy Comparison: The Warm-Start Experiment
A natural question arises: if MCMC refinement is needed for $n=4$, why use gradient descent at all? We conducted a controlled experiment comparing four initialization strategies for MCMC, measuring convergence speed (steps to reach 99.99% accuracy):
Experimental Setup.
For each of the 10 operations, we run MCMC refinement from four initialization conditions with 3 random seeds each:
  • (A) Random ternary: Initialize mask uniformly from $\{-1,0,+1\}^{16}$
  • (B) WHT-threshold: Quantize exact Walsh-Hadamard coefficients (threshold = 0.1)
  • (C) GD warm-start: Train Gumbel-Softmax model for 2000 steps from random init, then quantize
  • (D) WHT→GD: Initialize GD with WHT-threshold mask, train 2000 steps, then quantize
Results: WHT Dominates, GD Learns Topology Without Helping Convergence.
The experiment reveals a surprising result: WHT-threshold initialization achieves 100% accuracy in 110 ± 287 MCMC steps, dramatically outperforming all other methods. Random initialization requires $1387\pm 706$ steps, while GD warm-start actually underperforms random at $1557\pm 557$ steps.
However, GD is not futile: tracking Jaccard similarity during warm-up reveals that GD achieves AUC(Jaccard) = 0.732, confirming it learns the support topology. The puzzle resolves when recognizing that soft Gumbel-Softmax weights don’t quantize well for discrete MCMC—the learned topology is present in the weight magnitudes, but the hard quantization destroys the fine-grained balance GD discovered. WHT, computing exact Fourier coefficients directly from the truth table, provides initialization that MCMC can immediately refine.
Implications.
This validates the spectral synthesis pipeline: for $n\geq 4$, exact coefficient computation via WHT combined with ternary quantization provides superior initialization compared to gradient-based warm-starting. The “topology first, definition second” principle holds—but for discrete synthesis at scale, topology discovery via exact spectral analysis beats gradient descent.
@@ 7.5 Initialization Strategy Comparison: The Warm-Start Experiment
7.6 Unified Diagnostic Summary
Figure 10 synthesizes the diagnostic results across all four phases, validating the “topology first, definition second” thesis. Panel A shows Jaccard AUC progression: near-perfect at Phase 1 ($n=2$), declining but non-zero at Phase 3 ($n=3$) where accuracy plateaus, and moderate at Phase 4 ($n=4$) with GD warm-start. Panel B confirms high final accuracy across Phases 1-3. Panel C demonstrates the warm-start experiment result: WHT initialization (110 steps) dramatically outperforms random (1387 steps), pure GD (1557 steps), and WHT→GD (1393 steps). Panel D summarizes key findings, showing the transition from pure gradient learning at $n=2$ to hybrid spectral synthesis at $n\geq 3$.
Figure 10: Unified Diagnostic Summary Across All Phases. (A) AUC(Jaccard) comparison showing topology learning quality across phases. (B) Final accuracy across Phases 1-3. (C) Phase 4 warm-start MCMC convergence: WHT initialization achieves fastest convergence. (D) Key insights summary validating “topology first, definition second” thesis and showing where exact spectral methods (WHT) surpass gradient-based approaches.

### v3

@@ Abstract
Learning precise Boolean logic via gradient descent remains challenging: neural networks typically converge to “fuzzy” approximations that degrade under quantization. We introduce Hierarchical Spectral Composition, a transparent-by-design architecture that selects spectral coefficients from an interpretable Boolean Fourier basis and composes them via Sinkhorn-constrained routing with column-sign modulation. Each Fourier coefficient $\hat{f}(S)$ has explicit semantic meaning—the correlation of $f$ with parity of variables SS—making learned representations intrinsically explainable. Our approach draws on Manifold-Constrained Hyper-Connections (mHC) [1], adapting Birkhoff polytope projection from LLM training stability to logic synthesis, with column-sign modulation enabling Boolean negation.
We validate across five phases: (1–3) For $n=2,3,4$, we achieve 100% accuracy on canonical Boolean operations via gradient descent ($n=2$), exhaustive enumeration ($n=3$), and spectral synthesis with MCMC refinement ($n=4$), all converging to ternary masks $\{-1,0,+1\}$ enabling single-cycle GPU inference at 10,959 MOps/s. (4) Exact Walsh-Hadamard transforms scale to $n=28$ (268M coefficients, 1.64B coeffs/sec). (5) Oracle learning experiments ($n=16$) compare five coefficient estimation methods (Monte Carlo, Goldreich-Levin, spectral filtering) on parity, majority, and comparator functions, revealing that symbolic structure beats black-box learning: exploiting symmetry constraints boosts majority accuracy from 72% (MC) to 86% (GL+Symmetry), a +38% gain ($p<0.001$), while Birkhoff projection aids routing but not coefficient denoising. These findings establish that domain knowledge integration—encoding known function properties as hard spectral constraints—produces both higher accuracy and greater interpretability than generic learning algorithms, aligning with the vision of explainable neurosymbolic AI.
@@ 1 Introduction
Explainability and Transparent-by-Design AI.
A central motivation for spectral methods is interpretability: each Fourier coefficient $\hat{f}(S)$ has explicit semantic meaning as the correlation between $f$ and the parity character $\chi_{S}$. Unlike deep network activations that resist human interpretation, spectral coefficients are ground truth features defined by classical Boolean function analysis [3]. Our architecture’s transparency extends beyond coefficient interpretability: (1) ternary weights $\{-1,0,+1\}$ are human-readable (no hidden floating-point parameters), (2) Sinkhorn routing matrices are column-stochastic (each child is a convex combination of parents, preserving interpretability through composition), and (3) learned logic can be verified via formal methods or exhaustive testing (unlike black-box neural approximations). The oracle learning experiments (Phase 5) further demonstrate that symbolic knowledge integration—encoding known structural properties (symmetry, degree bounds) as hard constraints on the Fourier basis—yields both higher accuracy and greater explainability than generic learning algorithms. This aligns with the vision of explainable neurosymbolic AI: combining continuous optimization with discrete symbolic structures to produce models whose reasoning processes are open to inspection and human-level understanding.
@@
Our contributions span architecture design, theoretical analysis, and empirical validation across five phases:
@@
    • Scalability: Exact FWHT at 1.64B coeffs/sec ($n\leq 28$, 268M coefficients); hierarchical composition for 64-bit adders
@@
    Oracle Learning and Symbolic Knowledge Integration: We implement and compare five coefficient estimation methods (MC, Goldreich-Levin, spectral filtering) on three function families at $n=16$, demonstrating that symbolic structure beats black-box learning: exploiting symmetry constraints boosts majority accuracy from 72% to 86% (+38%, $p<0.001$), while establishing that Birkhoff projection aids routing but not coefficient denoising.
  6. 6.
@@ 2.1 Manifold-Constrained Routing: The mHC Foundation
2.2 Recent Differentiable Logic Gate Networks (2024-2025)
Three concurrent works directly compete with our spectral approach:
WARP-LUTs
(arXiv 2510.15655, 2025) represents our closest competitor, using Walsh-Hadamard spectral representation for differentiable LUT training via probabilistic relaxation. Critical distinction: WARP-LUTs employs soft probabilistic gates while our approach uses deterministic Sinkhorn-constrained routing with column-sign modulation. Our Proposition 3 proves doubly stochastic matrices cannot represent Boolean negation—a fundamental expressivity gap WARP-LUTs inherits from probabilistic methods.
Mind the Gap
[…]
Convolutional DLGNs
(Petersen et al., NeurIPS 2024 Oral) scales differentiable logic gate networks to 86.29% on CIFAR-10 using 61M logic gates via softmax relaxation over 16 gate types. Unlike their gate enumeration, we operate in the mathematically dual Fourier space—selecting spectral coefficients rather than enumerating gates. This duality offers complementary strengths: CDLGNs scale via convolution, we scale via spectral hierarchy.
@@ 2.3 Spectral Learning of Boolean Functions
The Fourier analysis of Boolean functions is foundational to computational learning theory [4, 5]. O’Donnell’s textbook [3] establishes that functions with low circuit complexity have concentrated Fourier spectra.
Recent Theoretical Advances.
Hidden Progress in Deep Learning (Barak et al., NeurIPS 2022) shows neural networks learn k-sparse parities via a Fourier gap mechanism: SGD gradually amplifies sparse spectral solutions, with progress invisible until phase transition. This mechanism relates to our explicit coefficient selection—we directly identify what SGD implicitly amplifies. Hardness of Learning Fixed Parities (Shoshani & Shamir, arXiv 2501.00817, January 2025) proves gradient descent on one-hidden-layer ReLU networks fails for any fixed parity, establishing new bounds on Fourier coefficient decay. Our explicit spectral coefficient selection potentially circumvents this hardness.
Spectral Bias and Regularization.
A Scalable Walsh-Hadamard Regularizer (Gorji et al., UAI 2023) proves neural networks are biased toward low-degree Walsh-Hadamard coefficients, hurting Boolean function generalization. Their solution: functional regularization to learn higher-degree coefficients. We invert this paradigm—directly selecting coefficients from a frozen spectral basis rather than regularizing against bias. The Frequency Principle (Xu et al., updated 2024) establishes DNNs fit target functions from low to high frequencies, explaining why gradient descent struggles with high-frequency Boolean functions. This foundational result justifies why domain-specific spectral structure outperforms black-box learning.
Spectral Methods for Interpretability.
Activation Spectroscopy (arXiv 2501.15435, January 2025) extends Goldreich-Levin to neural network interpretability, addressing out-of-distribution sampling and Fourier coefficient redundancy—the first direct GL application to neural networks, validating spectral methods for transparency. FourierSAT (Kyrillidis et al., AAAI 2020) uses Walsh-Fourier transforms to reduce SAT to continuous optimization. Critical distinction: FourierSAT is a solver for given formulas; our approach is a learner that discovers Boolean functions from data via gradient descent.
Our Contribution.
Unlike prior work applying Walsh-Hadamard compression [32] or parity learning [31], we embed the spectral basis as frozen primitives and learn to select and compose them via mHC-style routing, achieving interpretable coefficient representations with zero-loss quantization.
@@ 2.4 Neural Logic and Neuro-Symbolic Systems
2.5 Transparent-by-Design Architectures
The X-NeSy vision emphasizes models with built-in transparency rather than post-hoc explanations. Recent work validates this paradigm:
Interpretable Features.
ExplaiNN (Novakovsky et al., Genome Biology 2023) combines CNN expressiveness with linear model interpretability via Neural Additive Models, computing predictions as linear combinations from independent feature-specific models—producing interpretable learned coefficients analogous to our spectral coefficients. Shallow-ProtoPNet (Singh et al., Scientific Reports 2025) achieves full transparency using only one convolutional layer with no black-box components. tiSFM (Balcı et al., Bioinformatics/ISMB 2023) demonstrates domain-specific constraints (DNA sequence interpretation) enable inherently interpretable architectures without sacrificing performance.
Concept Bottleneck Models.
Energy-Based CBMs (Xu et al., ICLR 2024) unify prediction and concept intervention via energy-based frameworks. Stochastic CBMs (Vandenhirtz et al., NeurIPS 2024) model concept dependencies via multivariate normal distributions. Our spectral coefficients offer a mathematically grounded alternative: the Fourier basis provides complete, orthogonal feature representation rather than learned concepts requiring manual specification.
Mechanistic Interpretability.
Scaling Monosemanticity (Anthropic, May 2024) decomposes Claude 3 Sonnet into millions of interpretable features via sparse autoencoders. Transcoders (NeurIPS 2024) approximate MLP layers while maintaining input/weight contribution separation. Both demonstrate decomposition into interpretable features at scale—but both are post-hoc. Our transparent-by-design approach offers inherent interpretability through architecture: each Fourier coefficient $\hat{f}(S)$ has explicit semantic meaning (correlation with parity $\chi_{S}$), ternary weights are human-readable, and learned logic can be formally verified.
Our Contribution.
We demonstrate that Boolean Fourier analysis provides a mathematically principled foundation for transparent-by-design AI: spectral coefficients are ground-truth interpretable features defined by classical analysis [3], not learned representations requiring post-hoc interpretation.
@@ 3.2 The Birkhoff Polytope and Sinkhorn Projection
Proposition 2 (Sinkhorn Convergence to Permutation).
Let $P(\tau)=\mathrm{Sinkhorn}(\alpha,\tau)$ for fixed logits $\alpha\in\mathbb{R}^{m\times n}$ and temperature $\tau>0$. As $\tau\to 0$:
  1. (i)
    $P(\tau)$ converges to a permutation matrix $P^{*}\in\{0,1\}^{m\times n}$;
  2. (ii)
    The rate is controlled by the spectral gap:
    $\|P(\tau)-P^{*}\|_{F}=O\!\left(e^{-\Delta^{*}/\tau}\right)$ (6)
    where $\Delta^{*}$ is the gap between the optimal and second-best assignment costs;
  3. (iii)
    Zero-loss quantization: if $P(\tau)$ achieves $100\%$ classification accuracy for any $\tau>0$, then $P^{*}=\lim_{\tau\to 0}P(\tau)$ also achieves $100\%$ accuracy.
Proof sketch.
Part (i) follows from the theory of entropy-regularized optimal transport: as $\tau\to 0$, the entropic regularizer vanishes and the Sinkhorn output converges to the solution of the unregularized assignment problem [22]. Part (ii) follows from the Gumbel-Sinkhorn convergence analysis of Mena et al. [24], where the exponential rate $e^{-\Delta^{*}/\tau}$ arises from the Boltzmann distribution concentrating on the optimal permutation. Part (iii) follows from continuity: the routing function $x\mapsto\text{sign}(P\cdot x)$ is constant on connected components where $P\cdot x\neq 0$; since $P(\tau)\to P^{*}$ and accuracy is a discrete-valued function of routing, 100% accuracy is preserved in the limit. ∎
This proposition provides a theoretical foundation for the empirical “zero-loss quantization” observed in Phases 2–4: the Sinkhorn temperature anneal is not merely a heuristic but a provably convergent procedure. The spectral gap $\Delta^{*}$ in part (ii) is precisely the quantity measured in our spectral trajectory analysis (§9.4.1), where Sinkhorn’s monotonically decreasing $\Delta$ trajectory confirms controlled convergence while Gumbel-STE’s oscillating $\Delta$ indicates violation of this condition.
@@ 6.2 Representability Analysis
Beyond the 10 target operations, we exhaustively verify that every Boolean function on 3 variables admits a ternary PTF:
Theorem 2 (Universal Ternary Representability for $n=3$).
Every Boolean function $f:\{-1,+1\}^{3}\to\{-1,+1\}$ can be expressed as:
$f(x)=\text{sign}\!\Bigl(\sum_{S\subseteq\{1,2,3\}}w_{S}\,\chi_{S}(x)\Bigr),\quad w_{S}\in\{-1,0,+1\}$ (20)
Proof by exhaustive verification.
There are $|\{-1,+1\}^{\{-1,+1\}^{3}}|=2^{2^{3}}=256$ Boolean functions on 3 variables and $3^{2^{3}}=6{,}561$ ternary weight vectors in $\{-1,0,+1\}^{8}$. For each function, we evaluate all 6,5616{,}561 candidate masks against the full truth table. Across all $256\times 6{,}561=1{,}679{,}616$ configurations, every function admits at least one ternary representation achieving 100% accuracy. ∎
The support distribution (number of nonzero weights) among minimal-support witnesses reveals the structure of the ternary PTF landscape: mean support $5.1/8$, with the mode at support $6$ (38.3% of functions) and minimum support $1$ (the two constant functions).
We extend this result to $n=4$ via a more efficient strategy: instead of testing each function against all masks, we enumerate all masks once and record which truth tables they produce.
Theorem 3 (Universal Ternary Representability for $n=4$).
Every Boolean function $f:\{-1,+1\}^{4}\to\{-1,+1\}$ can be expressed as:
$f(x)=\text{sign}\!\Bigl(\sum_{S\subseteq\{1,2,3,4\}}w_{S}\,\chi_{S}(x)\Bigr),\quad w_{S}\in\{-1,0,+1\}$ (21)
Proof by exhaustive verification.
There are $2^{2^{4}}=65{,}536$ Boolean functions on 4 variables and $3^{2^{4}}=43{,}046{,}721$ ternary weight vectors in $\{-1,0,+1\}^{16}$. We enumerate all masks in batches, computing $\text{sign}(H\cdot w)$ where $H$ is the $16\times 16$ Fourier basis matrix, and recording each resulting truth table as a 16-bit integer. Full coverage of all 65,53665{,}536 truth tables is achieved after scanning $\approx 36\times 10^{6}$ of the $43\times 10^{6}$ masks. ∎
The $n=4$ support distribution has mean $10.8/16$, with the mode at support $10$ ($26.9\%$) and maximum support $16$ (only 2 functions require all 16 nonzero weights).
NPN Equivalence Classes.
Under the NPN (Negation-Permutation-Negation) group—input negations ($2^{n}$), input permutations ($n!$), and output negation (22)—the $2^{2^{n}}$ functions collapse into far fewer equivalence classes:
  • $n=3$: 256 functions $\to$ 14 NPN classes (group size $|G|=96$)
  • $n=4$: 65,536 functions $\to$ 222 NPN classes (group size $|G|=768$)
Since Theorems 23 establish representability for all functions, every NPN class is representable. This symmetry reduction is relevant for scaling: at $n=5$ ($2^{32}\approx 4\times 10^{9}$ functions), exhaustive enumeration is infeasible, but there are only 616,126616{,}126 NPN classes (OEIS A000370), making representative-based verification a viable path.
Together with Theorem 1 ($n=2$: 16/16), these results motivate:
Conjecture 1 (Universal Ternary Representability).
For all $n\geq 1$, every Boolean function $f:\{-1,+1\}^{n}\to\{-1,+1\}$ admits a ternary polynomial threshold function:
$f(x)=\text{sign}\!\Bigl(\sum_{S\subseteq[n]}w_{S}\,\chi_{S}(x)\Bigr),\quad w_{S}\in\{-1,0,+1\}$ (22)
Remark 4 (Connection to Fourier $L^{1}$ norms).
[…]
Heuristic Search at $n=5$ and $n=6$.
For $n\geq 5$, exhaustive verification is infeasible ($3^{32}\approx 1.9\times 10^{15}$ for $n=5$; $3^{64}\approx 3.4\times 10^{30}$ for $n=6$). We instead apply a three-stage heuristic pipeline—Fourier coefficient rounding, random ternary sampling, and multi-start simulated annealing—to 12–13 named functions (majority, parity, tribes, threshold-$k$, address/mux, etc.) and 500 ($n=5$) or 100 ($n=6$) random Boolean functions.
Table 6: Counterexample search results. “Named” = structured functions (majority, parity, threshold-$k$, tribes, etc.). “Random” = uniformly random truth tables. Failures indicate search limitations, not proven non-representability.
$n$ dim Named Random % Found Status
2 4 16/16 100% Proved (Thm 1)
3 8 256/256 100% Proved (Thm 2)
4 16 65,536/65,536 100% Proved (Thm 3)
5 32 12/12 397/500 79.4% Heuristic
6 64 12/13 41/100 41% Heuristic
At $n=5$, all named functions are representable via Fourier rounding alone (support ranging from $1$ for parity to $32$ for AND/OR). For random functions, 57.6% are found by Fourier rounding, 11.6% by simulated annealing in the initial pass, and 10.2% by extended SA—yielding 79.4% total. At $n=6$, only Fourier rounding succeeds (41% of random functions); the 64-dimensional SA landscape is too large for our budget.
These “not found” results are search failures, not counterexamples: our heuristic samples a vanishing fraction of the $3^{2^{n}}$ mask space. The sharp drop from 100% (proved, $n\leq 4$) to 79% ($n=5$, heuristic) to 41% ($n=6$) reflects increasing search difficulty, not necessarily decreasing representability. Conjecture 1 remains open; resolving it for $n=5$ via the NPN reduction (616,126 class representatives rather than $2^{32}$ functions) is a natural next step.